OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5 A.y – 1 = 5(x + 9) B.y – 9 = 5(x + 1) C.y – 9 = 5(x – 1) D.y + 9 = 5(x – 1) All help is appreciated :)

OpenStudy (anonymous):

The answer is C.

OpenStudy (anonymous):

Just plug in the coordinates in the third equation: y - 9 = 5(x - 1) 9 - 9 = 5(1 - 1) 0 = 5(0) 0 = 0 So C is the correct answer.